OpenStudy (anonymous):

0.3x-0.2y=4 0.5x+0.3y= -7/17

OpenStudy (saifoo.khan):

where the heck are my medals? lol.

OpenStudy (saifoo.khan):

Jk.

OpenStudy (saifoo.khan):

x = 100/17 y = -190/17

OpenStudy (anonymous):

that answer is wrong

OpenStudy (saifoo.khan):

i used my Calculator man.. it cant be wrong.

OpenStudy (anonymous):

it is wrong

OpenStudy (anonymous):

If you rationalize the coefficients, then you obtain exact solutions. @saifoo.khan's solutions are correct.\[\left\{x\to \frac{100}{17},y\to -\frac{190}{17}\right\}= \{x\to 5.88235,y\to -11.1765\} \]

jhonyy9 (jhonyy9):

0,3x-0,2y=4 - after you multiplied every two equation with 10 you get 0,5x+0,3y=-7/17 3x-2y=40 ---- y=(40-3x)/(-2) 5x-3y=-70/17 ----- 85x-51y=-70 --- 85x-51(-40+3x)/2 =-70 ------- 85x+(2040-153x)/2 =-70 ---- 170x +2040 -153x=-140 ----- 17x=-2180 ----- x=(-2180)/17 - than x=-2180/17 --- y=(40-3(-2180/17))/(-2) y=(680+6540)/(-34) y=7220/(-34)=(-3610)/17 x=(-2180)/17 y=(-3610)/17

OpenStudy (anonymous):

@jhonyy9, You made an error in line four of your solution. For some reason you changed the sign of "+0,3y" to "-0,3y" Compare line 4 to line 2. Your solutions for x and y will not verify when plugged into the second equation of the problem statement.

jhonyy9 (jhonyy9):

thank you @robtobey - yes -sorry

OpenStudy (anonymous):

You had me going for a while.